OpenStudy (anonymous):

A plane bisects a 90degree dihedral angle. From a point on this plane 16 in. from the common edge, perpendicular lines are constructed to the respective faces of the dihedral angle. Find the length of each perpendicular. #HELP

OpenStudy (anonymous):

First, do you know what a dihedral angle is?

OpenStudy (anonymous):

it is the angle formed between two intersecting planes. Am I right?

OpenStudy (anonymous):

yesh you are ok, just a second let me read it over. the question is poorly worded, no wonder you are confused

OpenStudy (anonymous):

Thank you so much! :D

OpenStudy (anonymous):

My maths textbook explains the steps to this sort of, the points on the planes are the same as the bisected section. The points should be an equal distant (equidistant, i think) from all of the sides and the perpendicular sides form two congruent angles. It says that if the point is 16 in from the common edge, the distance of it from the sides will be figured out like this. \[16/\sqrt{2}\]\[8/\sqrt{2}\]\[11.3\]

OpenStudy (anonymous):

Excuse me if there are errors in my response, I really suck at maths. I try and practice by helping people sometimes.

OpenStudy (anonymous):

ughm, how did 16/√2 become 8/√2?

OpenStudy (anonymous):

just a moment, thats not what was supposed to be it was supposed to be \[8\sqrt{2}\]

OpenStudy (anonymous):

oops :P

OpenStudy (anonymous):

its supposed to be 8sqrt(2) not 8/sqrt(2)

OpenStudy (anonymous):

thanks for pointing out that error, the equation editor confuses me sometimes

OpenStudy (anonymous):

so how did 16√2 become 8√2? I'm a little bit confused :(

OpenStudy (anonymous):

no, thats not it. 16 DIVIDED BY [square root of 2] EQUALS 8 TIMES [square root 2]

OpenStudy (anonymous):

Simplified it to 8sqrt(2), if you check with a calculator, both equal the same thing 11.313708499

OpenStudy (anonymous):

You don't need to simplify it, i was just showing steps.

OpenStudy (anonymous):

ahhhh I get it thank you so much! :)

OpenStudy (anonymous):

You are welcome, please give me a medal if I helped (best answer button)

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

ughm can you answer my other question pls? :D

OpenStudy (anonymous):

ughm emoticon, how did the answer become 16/sqrt of 2 or 8sqrt of 2? @emoticon